OpenStudy (anonymous):

solve using the square root property: 4a^2-45=0

OpenStudy (anonymous):

Let, 4a2 = 45, a2 = 45/4, a = radical 45/4.....

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

a = +/- radical 45/4.....

OpenStudy (anonymous):

the answer is positive or negative 3 radical 5 divided by 2 i just dont know how to get it

jimthompson5910 (jim_thompson5910):

\[\Large 4a^2 - 45 = 0\] \[\Large 4a^2 = 45\] \[\Large a^2 = \frac{45}{4}\] \[\Large a = \pm\sqrt{\frac{45}{4}}\] \[\Large a = \pm\frac{\sqrt{45}}{\sqrt{4}}\] \[\Large a = \pm\frac{3\sqrt{5}}{2}\]

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!